OpenStudy (freemap):

What is the distance between R(4, –7) and S(10, 4)?

pooja195 (pooja195):

Use the distance formula \[\huge d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]

Directrix (directrix):

Use the distance formula.

OpenStudy (freemap):

I'm working it out just wanted to post the options

Directrix (directrix):

Okay.

pooja195 (pooja195):

\[\huge d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] We are given: x1=4 y1=-7 x2=10 y2=4 Simply plug those in and you'd get your answer . Be sure to show your work so we can identify where you made the mistake if you make one

Directrix (directrix):

Apply above to get:

pooja195 (pooja195):

^solve from there.

Directrix (directrix):

Crank that out and post what you get. @freemap

OpenStudy (freemap):

12.53

pooja195 (pooja195):

Don't square it

OpenStudy (freemap):

A

pooja195 (pooja195):

yep

pooja195 (pooja195):

Well done :-)

OpenStudy (freemap):

Thanks

Directrix (directrix):

Look at the options. Yes.

pooja195 (pooja195):

yw and ty directrix :)

OpenStudy (freemap):

Always thanks @Directrix